OpenStudy (snowcrystal):

Help please I don't understand this...I will fan and medal!!!!! What is a possible value for the missing term of the geometric sequence? 1250, ?, 50,..... A.) 1200 B.) 650 C.) 250 D.) 125

rishavraj (rishavraj):

a = 1250 ar = ?? ar^2 = 50

rishavraj (rishavraj):

a is the first term ar second term ar^2 third term

OpenStudy (snowcrystal):

okay

OpenStudy (snowcrystal):

@K12awesomeness can you help

OpenStudy (k12awesomeness):

Just a sec..

OpenStudy (k12awesomeness):

Let x be the missing number Then 1250/ x = x /50 x^2 = 1250 x 50 = 62500 x =250

OpenStudy (k12awesomeness):

x = square root of 62500 = 250

OpenStudy (k12awesomeness):

It's C...

OpenStudy (snowcrystal):

Thank you @K12awesomeness

OpenStudy (k12awesomeness):

No Problem!!